ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 51, 68 Askin, Suat, Moscow Pedagogical University, Russia Additional Computer Geometry courses for advanced students The development of technology demands permanent modernization of school mathematics. We propose additional courses for advanced students: “Computer Geometry”. It includes discussion of well-known geometrical problems mainly Steiner problem, Voronoj problem, elements of fractal theory and elements of percolation theory etc. We choose these problems because of:

1. . The conditions of these problems don‘t demand studying the special part of mathematics,

2. . Discussion of the problems permits to relate the notions of abstract and concrete,

3. . Students can see the role of interface and processor speed. ( Computer languages Pascal, C , Mathematica 2.2, Mathematica 3.0